Find the probability distribution of the sum of the numbers obtained when two dice are thrown once.

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
The sample space contains 6*6=36 elements
sum of elements on dice from 2,3,4,.,12
p(sum=2) =(1,1)= 1/36 ;p(sum=3) =(1,2)(2,1)= 2/36 ;
p(sum=4) = (3,1)(2,2)(1,3)=3/36
p(sum=5) = (1,4)(2,3)(3,2)(4,1)=4/36
;p(sum=6) = (1,5)(5,1)(2,4)(4,2)(3,3)=5/36
;p(sum=7) = (1,6)(6,1)(2,5)(5,2)(4,3)(3,4)=6/36
p(sum=8) = (2,6)(6,2)(5,3)(3,5)(4,4)=5/36
;p(sum=9) = (6,3)(3,6)(4,5)(5,4)=4/36
;p(sum=10) = (6,4)(4,6)(5,5)=3/36
p(sum=11) = (5,6)(6,5)=2/36 ;
p(sum=12) = (6,6)=1/36
Hence expectation=np
=2(1/36)+3(2/36)+4(3/36)+5(4/36)+6(5/36)+7(6/36)+8(5/36)+9(4/36)+10(3/36)+11(2/36)+12(1/36)
=7
Thanks and Regards
Shaik Aasif
askIITians faculty